# Ambient Noise Full Waveform Inversion with Neural Operators

### v2

@@ Abstract
Numerical simulations of seismic wave propagation are crucial for investigating velocity structures and improving seismic hazard assessment. However, standard methods such as finite difference or finite element are computationally expensive. Recent studies have shown that a new class of machine learning models, called neural operators, can solve the elastodynamic wave equation orders of magnitude faster than conventional methods. Full waveform inversion is a prime beneficiary of the accelerated simulations. Neural operators, as end-to-end differentiable operators, combined with automatic differentiation, provide an alternative approach to the adjoint-state method. Since neural operators do not involve the Born approximation, when used for full waveform inversion they have the potential to include additional phases and alleviate cycle-skipping problems present in traditional adjoint-state formulations. In this study, we demonstrate the first application of neural operators for full waveform inversion on a real seismic dataset, which consists of several nodal transects collected across the San Gabriel, Chino, and San Bernardino basins in the Los Angeles metropolitan area.
@@ 1 Introduction
Seismic tomography for sedimentary basins is important for assessing earthquake hazards considering the fact that the basins can trap and amplify seismic waves, leading to stronger and longer ground shaking. Fine-scale subsurface structures are typically determined by active source surveys, which often operate at high frequencies but are expensive and limited by environmental impact issues. Alternatively, seismic data from passive sources, such as ambient noise and earthquakes, can be more readily acquired and can also compensate for the absence of low-frequency information in active source data. It has been shown in seismic interferometry that the cross-correlation between diffuse waveforms recorded at two stations approximates the elastodynamic Green’s function (Wapenaar, 2004; Wapenaar et al., 2010). The so-called empirical Green’s function (EGF) created from ambient noise is more commonly used by ray theory-based tomography methods (Guo et al., 2015; Lin et al., 2008; Yao et al., 2010; Zheng et al., 2011), which do not account for the finite-frequency effects in heterogeneous media. Full waveform inversion (FWI), which accounts for the full physics of wave propagation, can reveal more accurate and detailed subsurface structures (Liu et al., 2017, 2023; Maguire et al., 2022; Sager et al., 2020; Wang et al., 2021).
@@
Despite its significant advantages, FWI can be time-consuming and memory-intensive for large problems, which hinders its broader applicability. Moreover, the adjoint-state method (Fichtner et al., 2006a; Liu & Tromp, 2006; Plessix, 2006; Tromp et al., 2005) is conventionally used for FWI, which typically lies within the framework of the Born approximation (Hudson et al., 1977; Miles, 1960; Wu & Aki, 1985). The Born approximation linearizes the relationship between the perturbation to the wavefield and the perturbation to the medium and thus neglects multiple scattering effects, requiring the initial model to be close enough to the true Earth to avoid being trapped in local minima (Gauthier et al., 1986; Fichtner et al., 2006b; Panning et al., 2009; Virieux & Operto, 2009). In ambient noise tomography (ANT), the adjoint method usually requires a model derived from surface wave dispersion analysis and ray-based approximations as a starting point (Chen et al., 2014; Liu et al., 2023; Maguire et al., 2022; Zhang et al., 2018). To tackle these challenges, recent studies have shown that neural operators (Li et al., 2020a, b; Azizzadenesheli et al., 2024) can solve the wave equations orders of magnitude faster than traditional numerical solvers (Yang et al., 2021, 2023; Zou et al., 2024). Using automatic differentiation (AD) (Baydin et al., 2018; Rumelhart et al., 1986) instead of the Born approximation, neural operators have the potential to alleviate, though not fully resolve, the cycle-skipping problem in FWI (Métivier et al., 2018; Pladys et al., 2021; Yao et al., 2019; Sambridge et al., 2022), making it possible to start with a simpler velocity model (e.g., a 1D model).
@@
The primary advantage of neural operators in solving partial differential equations (PDEs) over other prevailing machine learning approaches, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), is that they are trained to learn the solution operator for an entire family of PDEs instead of a specific instance. This means that a neural operator has the potential to generalize to arbitrary PDE coefficients (e.g., velocity structures) if the training instances are ideally sampled, which are commonly sampled from random fields for neural operators. The training is a one-time effort for the forward process and there is no further training for the neural operator in the inversion stage, where the velocity parameters are updated with gradients computed from AD. This is similar to traditional FWI but eliminates the need to generate the adjoint wavefield. This way, we monitor the data misfit and have the flexibility to add regularization as needed.
@@
Our contributions are as follows. We demonstrate the first application of neural operators for ambient noise tomography on a real seismic dataset, specifically the Basin Amplification Seismic Investigation (BASIN) survey (BASIN, 2018; Clayton et al., 2019), which deployed linear nodal arrays in the northern Los Angeles (LA) basins. We provide a trained neural operator that is applicable without retraining to any linear array in the study area, or potentially other problems with a comparable spatial footprint. With the trained neural operator, FWI can be performed two orders of magnitude faster than the conventional adjoint method. Our tomography results are consistent with previous studies in the same region (Ghose et al., 2023; Li et al., 2023; Villa et al., 2023; Zou & Clayton, 2024) without the strong dependence on priors. The proposed method can be scaled to 3D, provided that sufficient computing resources are available.
2 Adjoint-State Method with the Born Approximation
Seismic wave propagation in an elastic medium follows the elastic wave equation
where $\mathbf{u}$ is the displacement wavefield, $\mathbf{f}$ is the body force, $\rho$ is the density, $\lambda$ and $\mu$ are the Lamé parameters. Equation 1 can be restated concisely with an integro-differential operator $\mathcal{L}$ as
$\mathcal{L}\mathbf{u}=\mathbf{f}.$ (2)
In seismology the medium is commonly parametrized by P-wave velocity $V_{P}$ and S-wave velocity $V_{S}$:
It is assumed that a small perturbation to the medium will perturb the operator $\mathcal{L}_{0}$ by $\delta\mathcal{L}$ and the wavefield $\mathbf{u}_{0}$ by $\delta\mathbf{u}$. Applying Equation 2 to the perturbed medium results in
$(\mathcal{L}_{0}+\delta\mathcal{L})(\mathbf{u}_{0}+\delta\mathbf{u})=\mathbf{f}.$ (4)
Since $\mathcal{L}_{0}\mathbf{u}_{0}=\mathbf{f}$, this leads to
$\mathcal{L}_{0}\delta\mathbf{u}+\delta\mathcal{L}\mathbf{u}_{0}+\delta\mathcal{L}\delta\mathbf{u}=\mathbf{0}.$ (5)
Neglecting the higher-order term $\delta\mathcal{L}\delta\mathbf{u}$, the first-order Born approximation is defined as
$\mathcal{L}_{0}\delta\mathbf{u}=-\delta\mathcal{L}\mathbf{u}_{0}.$ (6)
This indicates that the perturbation to the wavefield can be computed by treating the right-hand side as the source. It also implies that the Born approximation only accommodates single scattering and that it will break down if the perturbation is large relative to the background medium (Panning et al., 2009). Here, FWI is a local optimization strategy, where the solution is sought in the vicinity of an initial model by minimizing an objective function $\Phi$:
$\Phi=\frac{1}{2}\left\|\mathbf{u}_{r}-\mathbf{d}\right\|_{2}^{2}\quad\text{where}\quad\mathbf{u}_{r}=\mathbf{R}\mathbf{u},$ (7)
where $\mathbf{u}_{r}$ is the simulated waveforms queried at receivers through a sampling operator $\mathbf{R}$, and $\mathbf{d}$ is the observed data. Based on the Born approximation, the gradients of the objective with respect to parameters of interest $\mathbf{m}$ can be estimated as the cross-correlation between the forward wavefield $\mathbf{u}$ and the adjoint wavefield $\mathbf{v}$ at zero time lag (Louboutin et al., 2019):
$\nabla_{\mathbf{m}}\Phi=\sum_{i=1}^{N_{t}}\mathbf{u}[i]\mathbf{v}_{tt}[i],$ (8)
where $N_{t}$ is the number of time steps and the subscript $tt$ denotes a second-order time derivative. Equation 8 defines the adjoint-state method, in which the adjoint wavefield $\mathbf{v}$ is generated by back-propagating the residual $\delta\mathbf{d}$ in the medium:
$\mathcal{L}\mathbf{v}=\mathbf{R}^{T}\delta\mathbf{d}\quad\text{where}\quad\delta\mathbf{d}=\mathbf{u}_{r}-\mathbf{d}.$ (9)
In summary, the adjoint-state method with the Born approximation requires the initial model to be sufficiently accurate, with a travel-time error of less than half the period (Beydoun & Tarantola, 1988; Virieux & Operto, 2009). Otherwise, the cycle-skipping phenomenon may arise, leading to convergence to a local minimum (Gauthier et al., 1986; Pladys et al., 2021).
@@ 3.1 Helmholtz Neural Operator with Automatic Differentiation
The Green’s function $G$ is the impulse response of a linear differential operator $\mathcal{L}$ defined on a bounded, open domain $D\subset\mathbb{R}^{d}$ with specified parameters $a$:
$\mathcal{L}G_{a}\left(x,s\right)=\delta\left(x-s\right)\quad\text{where}\quad x\in D,$ (10)
where $\delta$ is the Dirac delta function and $s$ is the location of a point source. The solution to Equation 2 can then be derived from
$u\left(x\right)=\int_{D}G_{a}\left(x,y\right)f\left(y\right)\mathrm{d}y,$ (11)
where $f$ refers to the source function. This describes a linear mapping from $f$ to $u$ provided that $a$ (e.g., velocity) is fixed. However, in seismic wave propagation, the direct mapping from $a$ to $u$ is nonlinear and unknown in closed form. Based on the universal approximation theorem, neural operators can approximate arbitrary nonlinear continuous operators (Hornik et al., 1989; Kovachki et al., 2023).
@@
The solution operator can be parameterized much more efficiently in memory when solving Equation 2 in the frequency domain (i.e., the Helmholtz equation) than in the time domain, for which the Helmholtz Neural Operator (HNO) was proposed by Zou et al. (2024). This is because different frequency components can be solved individually and, thus, parallelized. Solutions in the time and frequency domains can be combined via the (inverse) Fourier transform. In this study, we construct a U-shaped HNO comprising $L=8$ inner layers of Fourier Neural Operators (FNOs) (Li et al., 2020a) to approximate the solution operator for the elastic wave propagation:
$v_{0}\left(x\right)=\left(\mathcal{P}\,a\right)\left(x\right),$ (12)
@@
$v_{l+1}\left(x\right)=\sigma\left(\mathcal{W}_{l}\,v_{l}\left(x\right)+\mathcal{F}^{-1}\left(\mathcal{F}\left(\kappa_{l}\right)\cdot\mathcal{F}\left(v_{l}\right)\right)\right),$
$u\left(x\right)=\left(\mathcal{Q}\,v_{L}\right)\left(x\right),l=0,..,L-1,$
@@
where $a$ is the input including $V_{P}$, $V_{S}$, the source location, and a constant function indicating the frequency value, $u$ is the output wavefield solution, $v_{l}$ is input to the ${l+1}^{th}$ layer, $\mathcal{P}$ is a point-wise operator used to lift the dimension, $\mathcal{Q}$ is used to project the output to the desired dimension, $\mathcal{W}_{l}$ acts as a residual connection, $\kappa_{l}$ is a parametric kernel function, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and inverse Fourier transform, respectively, and $\sigma$ is a nonlinear activation function, which we set as GELU (Hendrycks & Gimpel, 2016). Figure 1 illustrates the model architecture and Table S1 provides the details in each FNO layer. We use FNOs to speed up the integral calculation, and the U-shape allows for a deeper model and skip connections (Li et al., 2020c; Rahman et al., 2022; Ronneberger et al., 2015).
@@
Figure 1: Model architecture. $a$ is the input including $V_{P}$, $V_{S}$, the source location, and a constant function indicating the frequency value. $u$ is the output wavefield solution. $\mathcal{P}$ is a point-wise operator used to lift the dimension. $\mathcal{Q}$ is used to project the output to the desired dimension. $\mathcal{B}$ is the inner integral operator chosen as the FNO, in which $v$ is input to the layer, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and inverse Fourier transform, respectively, $\mathcal{R}$ is a linear operator, $\mathcal{W}$ acts as a residual connection, and $\sigma$ is a nonlinear activation function. $\Lambda$ is a graph neural operator (GNO) used to query the waveforms at the free surface, which is parameterized as a neural network with $(96,96,96)$ neurons.
In the previous section, we showed that in the conventional adjoint method, the gradients for updating velocity parameters are computed based on the Born approximation. However, with neural operators (or other machine learning methods developed in the PyTorch framework), the gradients of the objective to any parameters in the computational graph are automatically computed using the chain rule (Rumelhart et al., 1986). For example, if the scalar-valued objective function $J$ is defined as
$J=\frac{1}{2}\sum_{i=1}^{N}\left(\hat{u}_{i}-u_{i}\right)^{2},$ (13)
where $N$ is the number of observations, $\hat{u}_{i}$ is the model prediction, and $u_{i}$ is the ground truth, the gradient of $J$ with respect to a parameter $m_{j}$ is given by
$\frac{\partial J}{\partial m_{j}}=\sum_{i=1}^{N}\frac{\partial J}{\partial\left(\hat{u}_{i}-u_{i}\right)}\frac{\partial\left(\hat{u}_{i}-u_{i}\right)}{\partial\hat{u}_{i}}\frac{\partial\hat{u}_{i}}{\partial m_{j}}=\sum_{i=1}^{N}\left(\hat{u}_{i}-u_{i}\right)\frac{\partial\hat{u}_{i}}{\partial m_{j}},$ (14)
where $\frac{\partial\hat{u}_{i}}{\partial m_{j}}$ is further derived through a chain of differential operations. This is termed the reverse-mode AD (Baydin et al., 2018; Elliott, 2018). Because all the functions, expressions, and control flow structures used in neural operators are differentiable and compatible with AD, it is unnecessary to run an adjoint simulation for gradient estimation. Furthermore, the linearization taken by the Born approximation is not made here, which will ease the need for close initial models. Note that this does not mean the HNO-AD method is independent of initial models, but rather that it only reduces the dependence. It is still likely to get stuck in local minima due to bad initialization.

### v3

@@ Abstract
Numerical simulations of seismic wave propagation are crucial for investigating velocity structures and improving seismic hazard assessment. However, standard methods such as finite difference or finite element are computationally expensive. Recent studies have shown that a new class of machine learning models, called neural operators, can solve the elastodynamic wave equation orders of magnitude faster than conventional methods. Full waveform inversion is a prime beneficiary of the accelerated simulations. Neural operators, as end-to-end differentiable operators, combined with automatic differentiation, provide an alternative approach to the adjoint-state method. State-of-the-art optimization techniques built into PyTorch provide neural operators with greater flexibility to improve the optimization dynamics of full waveform inversion, thereby mitigating cycle-skipping problems. In this study, we demonstrate the first application of neural operators for full waveform inversion on a real seismic dataset, which consists of several nodal transects collected across the San Gabriel, Chino, and San Bernardino basins in the Los Angeles metropolitan area.
@@ 1 Introduction
Seismic tomography for sedimentary basins is important for assessing earthquake hazards considering the fact that the basins can trap and amplify seismic waves, leading to stronger and longer ground shaking. Fine-scale subsurface structures are typically determined by active source surveys, which often operate at high frequencies but are expensive and limited by environmental impact issues. Alternatively, seismic data from passive sources, such as ambient noise and earthquakes, can be more readily acquired and can also compensate for the absence of low-frequency information in active source data. Many have shown in seismic interferometry that the cross-correlation between diffuse waveforms recorded at two stations approximates the elastodynamic Green’s function (Wapenaar, 2004; Wapenaar et al., 2010). The so-called empirical Green’s function (EGF) created from ambient noise is more commonly used by ray theory-based tomography methods (Guo et al., 2015; Lin et al., 2008; Yao et al., 2010; Zheng et al., 2011), which do not account for the finite-frequency effects in heterogeneous media. Full waveform inversion (FWI), which accounts for the full physics of wave propagation, can reveal more accurate and detailed subsurface structures (Liu et al., 2017, 2023a; Maguire et al., 2022; Sager et al., 2020; Wang et al., 2021a; Zhang et al., 2021).
@@
Despite its significant advantages, FWI can be time-consuming and memory-intensive for large problems, which hinders its broader applicability. Moreover, the adjoint-state method (Fichtner et al., 2006a; Liu & Tromp, 2006; Plessix, 2006; Tromp et al., 2005), which is conventionally used for FWI, requires the initial model to be sufficiently close to the true Earth to avoid being trapped in local minima (Gauthier et al., 1986; Fichtner et al., 2006b; Panning et al., 2009; Virieux & Operto, 2009). In ambient noise tomography (ANT), the adjoint-state method typically relies on a model derived from surface wave dispersion analysis and ray-based approximations as a starting point (Chen et al., 2014; Liu et al., 2023a; Maguire et al., 2022; Zhang et al., 2018, 2020). To tackle these challenges, recent studies have shown that neural operators (Li et al., 2020a, b; Azizzadenesheli et al., 2024) can solve the wave equations orders of magnitude faster than traditional numerical solvers (Yang et al., 2021, 2023; Zou et al., 2024). In deep learning, an analogue to the adjoint-state method is known as the reverse-mode automatic differentiation (AD) (Baydin et al., 2018; Rumelhart et al., 1986). While these two methods have been shown to be mathematically equivalent (Richardson, 2018; Wang et al., 2021b; Zhu et al., 2021), AD, which underpins the modern deep learning platform PyTorch (Paszke et al., 2019), remains aligned with advances in state-of-the-art optimization techniques. For example, methods such as mini-batching (LeCun et al., 2002), stochastic gradient descent (SGD) (Bottou, 2010), and adaptive moment estimation (Adam) (Kingma & Ba, 2014) have been developed to better navigate non-convex loss landscapes and escape poor local minima. These methods have prevailed for over a decade in the machine learning community, while remaining relatively underexplored in the classic FWI world. Only in recent years have they begun to attract attention (Bernal-Romero & Iturrarán-Viveros, 2021; Mao et al., 2025; Richardson, 2018; Sun et al., 2020; Thrastarson et al., 2022; van Herwaarden et al., 2020). Therefore, in alignment with the rapid advancements in optimization techniques, neural operators have the potential to alleviate — though not fully resolve — the cycle-skipping problem in FWI, making it possible to start from a simpler velocity model (e.g., a 1D model).
@@
The primary advantage of neural operators in solving partial differential equations (PDEs) over other prevailing machine learning approaches, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), is that they are trained to learn the solution operator for an entire family of PDEs instead of a specific instance. This means that a neural operator has the potential to generalize to arbitrary PDE coefficients (e.g., velocity structures) if the training instances are ideally sampled — for example, from random fields (Yang et al., 2021, 2023; Zou et al., 2024). The training is a one-time effort for the forward process and there is no further training for the neural operator in the inversion stage, where the velocity parameters are updated with gradients computed from AD. This is similar to traditional FWI but replaces the adjoint simulation with back-propagation (LeCun et al., 1988).
In recent years, neural operators have been increasingly applied to seismic wave propagation modeling. They have been used for modeling acoustic (Yang et al., 2021), elastic (Li et al., 2025; Yang et al., 2023), and viscoelastic (Wei & Fu, 2022) waves. Most applications are limited to 2D due to computational and memory constraints. However, Lehmann et al. (2024, 2025) circumvented the 3D computational bottleneck by predicting only the surface ground motion, while Zou et al. (2024) modeled the full 3D wavefield by parameterizing the operator in the frequency domain, thereby eliminating the need to model the time dimension. Learning frequency-domain solutions offers memory efficiency for neural operators and has attracted growing interest (Cheng et al., 2025; Huang & Alkhalifah, 2025; Kong et al., 2025; Li et al., 2023a; Zhang et al., 2023). Among these, Huang & Alkhalifah (2025) and Cheng et al. (2025) proposed learning the residual (scattered) wavefield instead of the full wavefield to address the point-source singularity. Current research has primarily focused on synthetic data, while the application of neural operators to real seismic data remains largely unexplored.
@@
Our contributions are as follows. We demonstrate the first application of neural operators for ambient noise tomography on a real seismic dataset, specifically the Basin Amplification Seismic Investigation (BASIN) survey (BASIN, 2018; Clayton et al., 2019), which deployed linear nodal arrays in the northern Los Angeles (LA) basins. We provide a trained neural operator that is applicable without retraining to any linear array in the study area, or potentially other problems with a comparable spatial footprint. With the trained neural operator, FWI can be performed two orders of magnitude faster than the conventional adjoint method. Our tomography results are consistent with previous studies in the same region (Ghose et al., 2023; Li et al., 2023b; Villa et al., 2023; Zou & Clayton, 2024) without the strong dependence on priors. The proposed method can be scaled to 3D, provided that sufficient computing resources are available.
@@ 2.1 Helmholtz Neural Operator with Automatic Differentiation
Seismic wave propagation in an elastic medium follows the (isotropic) elastic wave equation
where $\mathbf{u}$ is the displacement wavefield (three-component vector in 3D media), $t$ is time, $\nabla$ is the gradient with respect to the space coordinates, $\mathbf{f}$ is the body force (source term), $\rho$ is the density, $\lambda$ and $\mu$ are the Lamé parameters. Equation 1 can be restated concisely with an integro-differential operator $\mathcal{L}$ as
$\mathcal{L}\mathbf{u}=\mathbf{f}.$ (2)
In seismology the medium $\mathbf{m}$ is commonly parametrized by P-wave velocity $V_{P}$ and S-wave velocity $V_{S}$:
The direct mapping from physical parameters $\mathbf{m}$ to observed waveforms $\mathbf{d}$ is nonlinear and unknown in closed form:
$\mathbf{d}=\mathbf{G}(\mathbf{m}).$ (4)
@@
Based on the universal approximation theorem, neural operators can approximate arbitrary nonlinear continuous operators (Hornik et al., 1989; Kovachki et al., 2023). The solution operator can be parameterized much more efficiently in memory when solving Equation 2 in the frequency domain (i.e., the Helmholtz equation) than in the time domain (Huang & Alkhalifah, 2025; Kong et al., 2025; Zou et al., 2024). This is because different frequency components can be solved individually and, thus, parallelized. In this study, we employ the Helmholtz Neural Operator (HNO) proposed by Zou et al. (2024). The HNO computes the solution for each frequency component independently by defining the batch of data at the frequency level. In other words, a batch can contain multiple frequency components due to data shuffling. We use a single HNO to learn the entire frequency band of interest (0.10.10.1 to 0.50.50.5 Hz), although Kong et al. (2025) found that using multiple models to handle different frequencies slightly improves performance. Solutions in the time and frequency domains can be related via the (inverse) Fourier transform. We construct a U-shaped HNO comprising $L=8$ inner layers of Fourier Neural Operators (FNOs) (Li et al., 2020a), followed by a Graph Neural Operator (GNO) (Li et al., 2020b), to predict the waveforms at the free surface (data, denoted by $\mathbf{\hat{d}}$):
$\mathbf{v}_{0}\left(x\right)=\left(\mathcal{P}\,\mathbf{a}\right)\left(x\right),$ (5)
@@
$\mathbf{v}_{l+1}\left(x\right)=\sigma\left(\mathcal{W}_{l}\,\mathbf{v}_{l}\left(x\right)+\mathcal{F}^{-1}\left(\mathcal{F}\left(\kappa_{l}\right)\cdot\mathcal{F}\left(\mathbf{v}_{l}\right)\right)\left(x\right)\right)\quad\text{where}\quad l=0,.,L-1,$
$\mathbf{v}_{gno}\left(x\right)=\int_{B_{r}(x)}\kappa\left(x,y,\mathbf{v}_{L}\right)\mathbf{v}_{L}(y)\,\mathrm{d}y,$
$\mathbf{\hat{d}}\left(x\right)=\left(\mathcal{Q}\,\mathbf{v}_{gno}\right)\left(x\right),$
@@
where $\mathbf{a}$ is the input including $V_{P}$, $V_{S}$, the source location, and a constant function indicating the frequency value, $\mathbf{\hat{d}}$ is the predicted data for the given frequency component (including the real and imaginary parts), $\mathbf{v}_{l}$ is input to the ${l+1}^{th}$ FNO layer ($l=0,...,7$) or the GNO ($l=8$), $\mathbf{v}_{gno}$ is the GNO output, $\mathcal{P}$ is a point-wise operator used to lift the dimension, $\mathcal{Q}$ is used to project the output to the desired dimension, $\mathcal{W}_{l}$ acts as a residual connection, $\kappa_{l}$ is a parametric kernel function, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $\sigma$ is a nonlinear activation function, which we set as GELU (Hendrycks & Gimpel, 2016). The GNO uses a kernel function $\kappa$ parameterized as a three-layer neural network, which takes positional information and the output of the last FNO layer as input. $B_{r}$ denotes the predefined neighborhood of the queried position $x$, defined as all points below the corresponding surface location for computational efficiency. We use FNOs to speed up the integral calculation, and the U-shape allows for a deeper model and skip connections (Li et al., 2020c; Rahman et al., 2022; Ronneberger et al., 2015). The GNO can improve predictions at queried points and provide greater data efficiency (Zou et al., 2024). Figure 1 illustrates the model architecture and Table S1 provides the details in each neural operator layer.
@@
Figure 1: Model architecture. $\mathbf{a}$ is the input given by $V_{P}$, $V_{S}$, the source location, and a constant function indicating the frequency value. $\mathbf{\hat{d}}$ is the predicted data. $\mathcal{P}$ is a point-wise operator used to lift the dimension. $\mathcal{Q}$ is used to project the output to the desired dimension. $\mathcal{B}$ is the inner integral operator chosen as the FNO, in which $\mathbf{v}$ is input to the layer, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, $\mathcal{R}$ is a linear operator, $\mathcal{W}$ acts as a residual connection, and $\sigma$ is a nonlinear activation function. $\Lambda$ is a GNO used to query the waveforms at the free surface. Blue circles denote concatenation along the channel dimension.
In conventional FWI, the solution is obtained by minimizing a scalar-valued objective function $\Phi$, which can, for example, be defined as half the squared $\ell_{2}$-norm of the data residual:
$\Phi=\frac{1}{2}\left\|\mathbf{\hat{d}}-\mathbf{d}\right\|_{2}^{2}\quad\text{where}\quad\mathbf{\hat{d}}=\mathcal{S}\mathbf{u},$ (6)
where $\mathbf{\hat{d}}$ is the modeled (predicted) waveforms queried at receivers through a sampling operator $\mathcal{S}$ and $\mathbf{d}$ is the observed data . The parameters of interest $\mathbf{m}$ are updated iteratively using a gradient-based method. Using the adjoint-state method, the gradients of the objective with respect to parameters of interest can be estimated as the cross-correlation between the forward wavefield $\mathbf{u}$ and the adjoint wavefield $\mathbf{u}^{*}$ at zero time lag, weighted by an analytically derived term (Louboutin et al., 2019):
$\nabla_{\mathbf{m}}\Phi=-\mathbf{u}^{T}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{m}}\right)^{T}\mathbf{u}^{*}.$ (7)
The adjoint wavefield $\mathbf{u}^{*}$ is generated by back-propagating the residual $\delta\mathbf{d}$ in the medium:
$\mathcal{L}^{T}\mathbf{u}^{*}=\mathcal{S}^{T}\delta\mathbf{d}\quad\text{where}\quad\delta\mathbf{d}=\mathbf{\hat{d}}-\mathbf{d}.$ (8)
The adjoint-state method requires the initial model to be sufficiently accurate, with a travel-time error of less than half the period (Beydoun & Tarantola, 1988; Virieux & Operto, 2009; Alkhalifah, 2016). Otherwise, the cycle-skipping phenomenon may arise, leading to convergence to a local minimum (Gauthier et al., 1986; Pladys et al., 2021).
With neural operators (or other machine learning methods developed on the PyTorch platform), the gradients of the objective with respect to any parameters in the computational graph are automatically computed through back-propagation (Rumelhart et al., 1986). This is termed the reverse-mode AD (Baydin et al., 2018; Elliott, 2018). Provided that all the functions, expressions, and control flow structures used in neural operators are differentiable and compatible with AD, the adjoint simulation can be replaced by back-propagation (LeCun et al., 1988). It has been shown that the adjoint-state method and reverse-mode AD are mathematically equivalent (Richardson, 2018; Wang et al., 2021b; Zhu et al., 2021). However, the HNO-AD method eliminates the need to manually derive the gradient in Equation 7 on a case-by-case basis. Moreover, it can take advantage of state-of-the-art optimization techniques built into PyTorch — such as mini-batching and Adam — which help improve the optimization dynamics of FWI and mitigate cycle-skipping. Adam, in particular, a first-order gradient-based optimization algorithm that dynamically rescales the gradients for each parameter based on its past gradients and their squares (Kingma & Ba, 2014), has been shown to outperform other optimizers such as SGD and limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) (Bernal-Romero & Iturrarán-Viveros, 2021; Richardson, 2018; Sun et al., 2020).
We summarize the differences between the HNO-AD and adjoint-state methods as follows:
  1. 1.
    Forward modeling: In the adjoint-state method, forward modeling is performed by numerically solving the wave equation. In the HNO-AD method, it is performed via a mapping learned by a neural operator.
  2. 2.
    Gradient computation: In the adjoint-state method, the gradient is analytically derived and computed by running an adjoint simulation. In the HNO-AD method, this is replaced by automatic back-propagation.